1. At atmospheric pressure, a sample gas has a volume of 2.86 L at 29℃. What would its temperature be in Celsius if its volume increased to 2.86 L at the same pressure?



2.A balloon has a pressure of 2.05 atm at 28℃. If the temperature is reduced to 16℃, what will be the new pressure of the balloon?



3.A gas sample has a volume of 13.5 L at a pressure of 1.23 atm and a temperature of 34℃. What volume will the gas sample occupy at standard temperature and pressure (STP)? STP is 273.15 K and 1 atm, respectively.



4.Calculate the mass of 3.45 L SO2 at STP.



5.A gas sample is composed of 234 mol NO2 and 312 mol Cl2(g). determine the partial pressure of each gas if the sample’s total pressure is 30 atm.



1
Expert's answer
2021-11-22T08:42:01-0500

V1/T1= V2/T2

2.86/302=2.86/T2

T2= 302K =29°C


P1/T1=P2/V2

2.05/301=P2/289

0.068=P2/289

P2=289×0.068=1.97atm